### v1

# Persistent Homology via Finite Topological Spaces

Abstract
We propose a functorial framework for persistent homology based on finite topological spaces and their associated posets. Starting from a finite metric space, we associate a filtration of finite topologies whose structure maps are continuous identity maps. By passing functorially to posets and to simplicial complexes via crosscut constructions, we obtain persistence modules without requiring inclusion relations between the resulting complexes. We show that standard poset-level simplifications preserve persistent invariants and prove stability of the resulting persistence diagrams under perturbations of the input metric in a density-based instantiation.
2020 Mathematics Subject Classification. Primary: 55N31; Secondary: 06A11, 55U10.
Keywords: Persistent homology; finite topological spaces; posets; crosscut complexes; stability.
Bahçeşehir University, Faculty of Engineering
and Natural Sciences, Istanbul, Turkey
e-mail: selcuk.kayacan@bau.edu.tr
1 Introduction
Persistent homology has become a central tool in topological data analysis, providing multiscale descriptors of data through filtrations of simplicial complexes. Most standard constructions, such as Vietoris–Rips and Čech complexes, are built directly from metric data and rely on inclusion-based filtrations. While highly effective, these approaches often conceal order-theoretic and topological structures that arise naturally from the data and may be exploited prior to simplicialization.
In this paper, we develop a functorial framework for persistent homology that interpolates between metric data and simplicial complexes by way of finite topological spaces and posets. Starting from a finite metric space, we first associate to it a family of finite topologies indexed by a scale parameter. These topologies encode relational or density-based information and form a filtration via continuous identity maps. By passing functorially to associated posets and then to simplicial complexes via crosscut constructions, we obtain persistence modules without requiring inclusion relations between the resulting complexes.
This perspective separates the role of the metric as a generator of topological structure from the subsequent homological analysis. As a result, it allows for topological and combinatorial simplifications at the poset level, while preserving persistent invariants. Moreover, stability of the resulting persistence diagrams follows from the functoriality of the construction and elementary interleaving arguments.
Contributions
The main contributions of this paper are as follows:
  • We introduce a functorial pipeline from finite metric spaces to persistence modules via finite topological spaces and posets.
  • We show that the construction yields well-defined persistence modules even in the absence of inclusion relations between the associated simplicial complexes.
  • We employ crosscut complexes to pass from posets to simplicial complexes while preserving homotopy type.
  • We demonstrate that poset-level simplifications, such as beat-point removal, do not affect the resulting persistence diagrams.
  • We prove stability of the resulting persistence diagrams under perturbations of the input metric in a density-based instantiation.
2 Background
In this section, we shall review the basic definitions and results that are needed in this paper.
2.1 Persistent Homology
Persistent homology provides an algebraic framework for studying the evolution of topological features across a family of spaces indexed by a parameter. We briefly recall the definitions and results needed later; standard references include [4], [3], and [5].
[…]
Definition.
A persistence module indexed by $I$ is a covariant functor
[…]
Explicitly, this consists of a family of vector spaces $\{V_{i}\}_{i\in I}$ together with linear maps
$v_{i,j}:V_{i}\to V_{j}\quad\text{for all }i\leq j,$
such that $v_{i,i}=\mathrm{id}$ and $v_{j,k}\circ v_{i,j}=v_{i,k}$ whenever $i\leq j\leq k$.
Persistent homology is the study of persistence modules arising from filtrations via homology. Let $\mathbf{Top}$ denote the category of topological spaces and continuous maps.
Definition.
A filtration indexed by $I$ is a functor
$\mathbb{X}:I\longrightarrow\mathbf{Top}.$
[…]
Under suitable finiteness assumptions (e.g. pointwise finite-dimensionality), a persistence module admits a decomposition into interval modules, which can be encoded combinatorially.
Definition.
The persistence diagram $\mathsf{Dgm}(\mathbb{V})$ of a persistence module $\mathbb{V}$ is the multiset of points $(b,d)\in\overline{\mathbb{R}}^{2}$ corresponding to the birth and death parameters of its interval summands. Points on the diagonal $b=d$ are typically ignored, as they represent trivial features.
Two metrics play a central role in comparing persistence modules.
Definition.
The bottleneck distance $d_{B}$ between persistence diagrams is defined as the infimum over all matchings between diagrams of the maximum $L^{\infty}$-distance between matched points.
[…]
$\mathbb{W}(\varepsilon)(t):=\mathbb{W}(t+\varepsilon),$
with structure maps
$\mathbb{W}(\varepsilon)(t\leq s):=\mathbb{W}(t+\varepsilon\leq s+\varepsilon).$
Equivalently, $\mathbb{W}(\varepsilon)$ is obtained by precomposing $\mathbb{W}$ with the translation functor
$T_{\varepsilon}:(\mathbb{R},\leq)\longrightarrow(\mathbb{R},\leq),\qquad t\longmapsto t+\varepsilon.$
Intuitively, the shift $\mathbb{W}(\varepsilon)$ delays the appearance of all homological features by $\varepsilon$.
Definition.
Let $\varepsilon\geq 0$. Two persistence modules $\mathbb{V},\mathbb{W}$ indexed by $I\subset\mathbb{R}$ are said to be $\varepsilon$-interleaved if there exist natural transformations
$\mathbb{V}\Rightarrow\mathbb{W}(\varepsilon),\qquad\mathbb{W}\Rightarrow\mathbb{V}(\varepsilon),$
satisfying the standard commutativity conditions ensuring that compositions agree with the structure maps shifted by $2\varepsilon$. The interleaving distance $d_{I}(\mathbb{V},\mathbb{W})$ is the infimum of such $\varepsilon$.
A fundamental result in persistence theory states that the bottleneck distance $d_{B}$ and the interleaving distance $d_{I}$ coincide.
Theorem 1 (Isometry Theorem [3]).
For pointwise finite-dimensional persistence modules indexed by $\mathbb{R}$,
$d_{I}(\mathbb{V},\mathbb{W})=d_{B}(\mathsf{Dgm}(\mathbb{V}),\mathsf{Dgm}(\mathbb{W})).$
This theorem underpins the stability of persistent homology: small perturbations at the level of persistence modules lead to small changes in persistence diagrams. Finally, we recall a standard invariance principle.
Theorem 2 (Persistence Equivalence [4]).
[…]
This result allows one to replace a persistence module by any naturally isomorphic one without affecting its persistent invariants.
2.2 Finite Topological Spaces and Posets
Finite topological spaces admit a convenient combinatorial description in terms of preorders and posets. We briefly recall the basic constructions and results that will be used later; a comprehensive reference is [1].
Let $X$ be a finite topological space. For each $x\in X$, define
[…]
The collection $\{U_{x}\mid x\in X\}$ forms a minimal basis for the topology on $X$. As one can easily observe the relation $\leq$ on $X$ defined by
$x\leq y\quad\Longleftrightarrow\quad x\in U_{y}$
is a preorder. Conversely, given a finite preordered set $(X,\leq)$, one may recover a topology by declaring the sets
$U_{x}:=\{\,y\in X\mid y\leq x\,\},\qquad x\in X,$
to be a minimal basis. This establishes an equivalence between finite topological spaces and finite preorders.
Given a preorder $(X,\leq)$, define an equivalence relation by
[…]
The quotient $X/{\sim}$ inherits a natural topology whose associated preorder is a partial order.
Theorem 3 ([1, Proposition 1.3.1]).
Let $X$ be a finite topological space. Then $X$ is homotopy equivalent to the quotient space $X/{\sim}$.
In particular, every finite topological space is homotopy equivalent to a finite $T_{0}$-space, and finite $T_{0}$-spaces are in bijection with finite posets.
Let $\mathcal{P}$ be a finite poset.
Definition.
The order complex $\Delta(\mathcal{P})$ is the simplicial complex whose simplices are the finite chains in $\mathcal{P}$.
Order complexes provide a bridge between combinatorial and simplicial topology.
Theorem 4 ([1, Theorem 1.4.6]).
If $X$ is a finite $T_{0}$-space with associated poset $\mathcal{P}$, then $X$ is weakly homotopy equivalent to the order complex $\Delta(\mathcal{P})$.
Since $\Delta(\mathcal{P})$ captures the homotopy type of $X$ up to weak equivalence, the homology groups of $\Delta(\mathcal{P})$ and $X$ coincide.
While order complexes may be large, they often admit smaller simplicial models with the same homotopy type. Let $\mathcal{P}$ be a finite poset. A subset $\mathcal{C}\subseteq\mathcal{P}$ is called a crosscut if:
  1. (i)
    $\mathcal{C}$ is an antichain;
  2. (ii)
    every chain in $\mathcal{P}$ is comparable with at least one element of $\mathcal{C}$;
  3. (iii)
    for any $\mathcal{A}\subseteq\mathcal{C}$, if $\mathcal{A}$ has an upper bound or a lower bound in $\mathcal{P}$, then the join $\bigvee\mathcal{A}$ or the meet $\bigwedge\mathcal{A}$ exists in $\mathcal{P}$ (see [2] for details).
Definition.
Let $\mathcal{C}$ be a crosscut in a finite poset $\mathcal{P}$. The crosscut complex $\Delta(\mathcal{P},\mathcal{C})$ is the simplicial complex whose simplices are the subsets of $\mathcal{C}$ that admit a common upper bound or a common lower bound in $\mathcal{P}$.
Theorem 5 (Crosscut Theorem [6, Theorem 3], [2, Theorem 10.8]).
The crosscut complex $\Delta(\mathcal{P},\mathcal{C})$ is homotopy equivalent to the order complex $\Delta(\mathcal{P})$.
3 Construction
Let $(X,d)$ be a finite metric space. The metric on $X$ determines its pairwise distance matrix, which in turn encodes all geometric information about $X$ up to rigid motions and symmetries whenever $X$ arises as a finite subset of Euclidean space. This suggests that the distance matrix itself may be regarded as a primary data object, and that different constructions may extract topological information from this data at different scales or resolutions.
Our approach proceeds by associating to $(X,d)$ a family of topological spaces in a functorial manner, and then passing to combinatorial models suitable for homological analysis.
3.1 Topological Models from Metric Data
Any rule that assigns to $X$ a topology in a manner compatible with maps between metric spaces may be regarded as a topological model of the data. Given a topology $\mathcal{T}$ on $X$, we obtain a finite topological space $(X,\mathcal{T})$.
Associated to $(X,\mathcal{T})$ is its preorder defined by $x\leq y$ if and only if $x\in U_{y}$, where $U_{y}$ is the minimal basis element for the topology $\mathcal{T}$ containing $y$. Hence, one obtains a quotient space $X/{\sim}$ which is a finite $T_{0}$-space. The induced preorder on $X/{\sim}$ is in fact a partial order, yielding a finite poset $\mathcal{P}$.
From $\mathcal{P}$ we may form its order complex $\Delta(\mathcal{P})$, as well as crosscut complexes associated to suitable crosscuts.
Proposition 6.
Let $(X,\mathcal{T})$ be a finite topological space, $\mathcal{P}$ the associated poset, and $\mathcal{C}$ a crosscut consisting of maximal elements of $\mathcal{P}$. Then $(X,\mathcal{T})$ and the crosscut complex $\Delta(\mathcal{P},\mathcal{C})$ are weakly homotopy equivalent.
Proof.
By Theorem 3, $(X,\mathcal{T})$ is homotopy equivalent to its $T_{0}$-quotient. By Theorem 4, this quotient is weakly homotopy equivalent to the order complex $\Delta(\mathcal{P})$. The result now follows from the Crosscut Theorem (Theorem 5). ∎
Thus, any topology on $X$ determines a simplicial complex that is typically much smaller and more amenable to homological computation.
3.2 Filtrations and Functoriality
Let $I$ be a totally ordered finite index set. Suppose we are given a family of topologies
$\mathcal{T}_{1}\supseteq\mathcal{T}_{2}\supseteq\cdots\supseteq\mathcal{T}_{m}$
on the underlying set $X$. This defines a filtration
$\mathbb{X}\colon I\longrightarrow\mathbf{Top},$
by assigning $\mathbb{X}(i):=(X,\mathcal{T}_{i})$, and mapping each morphism $i\leq j$ in $I$ to the identity map $\mathrm{id}_{X}\colon(X,\mathcal{T}_{i})\longrightarrow(X,\mathcal{T}_{j})$, which is continuous since $\mathcal{T}_{j}\subseteq\mathcal{T}_{i}$.
Applying the constructions of Section 3.1 objectwise yields a corresponding filtration of simplicial complexes
$\mathbb{K}\colon I\longrightarrow\mathbf{Simp},\qquad i\longmapsto K_{i}:=\Delta(\mathcal{P}_{i},\mathcal{C}_{i}).$
Proposition 7.
For each $n\geq 0$, the persistence modules $H_{n}(\mathbb{X})$ and $H_{n}(\mathbb{K})$ are naturally isomorphic.
Proof.
Each step of the construction—passing from a finite topological space to its associated poset, and from a poset to a crosscut complex—is functorial with respect to continuous maps. Moreover, the weak homotopy equivalences involved are natural with respect to these maps. Applying homology yields naturally isomorphic persistence modules. In particular, the induced maps on homology commute with the structure maps of the filtrations. ∎
Remark.
Before forming the crosscut complexes, one may further simplify the posets $\mathcal{P}_{i}$ using standard techniques from poset combinatorics, such as removal of beat points or other homotopy-preserving reductions. Since these simplifications do not change the homotopy type, the resulting persistence diagrams remain unchanged by the Persistence Equivalence Theorem (Theorem 2).
Remark.
The filtration $\mathbb{K}$ does not consist of simplicial complexes related by inclusions, and therefore does not fit directly into the standard persistence homology pipeline based on matrix reduction. Nevertheless, more efficient algorithms than computing the homology of each $K_{i}$ independently are possible, exploiting the functorial structure of the construction. Developing such algorithms is an interesting direction for future work.
3.3 A Density-Based Filtration
We conclude this section by describing a concrete instance of the above framework. Let $t_{1}\leq t_{2}\leq\cdots\leq t_{m}$ be real parameters. For each $i$, let $\mathcal{O}_{i}$ be a collection of subsets of $X$ determined by a density criterion with threshold $t_{i}$, and let $\mathcal{T}_{i}$ be the topology generated by $\mathcal{O}_{i}$. The monotonicity of the density thresholds ensures that
$i\leq j\quad\Longrightarrow\quad\mathcal{T}_{j}\subseteq\mathcal{T}_{i},$
and hence defines a filtration as above.
In the extreme cases, $t_{1}$ may be chosen sufficiently small so that all co-singleton subsets are declared open, while $t_{m}$ may be chosen sufficiently large so that $\mathcal{T}_{m}$ is the trivial topology. Accordingly, the associated crosscut complexes interpolate between a discrete set of vertices and a single vertex.
Stability properties of this construction under perturbations of the metric data will be established in Section 4.
Remark.
Unlike Vietoris–Rips or Čech constructions, the simplicial complexes arising from our method are not designed to recover the ambient geometric shape of the data. In particular, a point cloud sampled from a circle need not give rise to a persistent one-dimensional homology class.
Instead, our construction summarizes statistical and relational information encoded in the density structure of the data. Persistent homology is used here as a multiscale descriptor of how dense regions interact and merge across thresholds, rather than as a tool for geometric shape reconstruction.
4 Stability
In this section we establish stability of the persistence modules arising from our construction under perturbations of the input metric. The proof relies on the functorial nature of the construction and on the theory of interleavings recalled in Section 2.1.
Let $(X,d)$ and $(X,d^{\prime})$ be two finite metric spaces on the same underlying set $X$. We say that $d^{\prime}$ is an $\varepsilon$-perturbation of $d$ if
$|d(x,y)-d^{\prime}(x,y)|\leq\varepsilon\quad\text{for all }x,y\in X.$
Such perturbations arise naturally, for example, from noise in measurements or discretization effects.
4.1 Stability of the Induced Filtrations
Fix a family of threshold values $\{t\}_{t\in\mathbb{R}}$ and let
$\mathbb{X}^{d},\mathbb{X}^{d^{\prime}}\colon(\mathbb{R},\leq)\longrightarrow\mathbf{Top}$
denote the filtrations obtained from $(X,d)$ and $(X,d^{\prime})$, respectively, via the density-based procedure described in Section 3.3.
Lemma 8.
If $d^{\prime}$ is an $\varepsilon$-perturbation of $d$, then for every $t\in\mathbb{R}$ there exist continuous identity maps
$(X,\mathcal{T}^{d}_{t})\longrightarrow(X,\mathcal{T}^{d^{\prime}}_{t+\varepsilon}),\qquad(X,\mathcal{T}^{d^{\prime}}_{t})\longrightarrow(X,\mathcal{T}^{d}_{t+\varepsilon}).$
Proof.
By construction, the topology $\mathcal{T}^{d}_{t}$ is generated by a family $\mathcal{O}^{d}_{t}$ of subsets of $X$, where membership of a subset $S\subseteq X$ in $\mathcal{O}^{d}_{t}$ depends only on whether its density (computed using $d$) exceeds the threshold $t$.
If $d^{\prime}$ is an $\varepsilon$-perturbation of $d$, then any density inequality defining membership in $\mathcal{O}^{d}_{t}$ with respect to $d$ remains valid at level $t+\varepsilon$ with respect to $d^{\prime}$. Consequently,
$\mathcal{O}^{d}_{t}\subseteq\mathcal{O}^{d^{\prime}}_{t+\varepsilon},\quad\text{and hence}\quad\mathcal{T}^{d^{\prime}}_{t+\varepsilon}\subseteq\mathcal{T}^{d}_{t}.$
This implies that the identity map on $X$ is continuous from $(X,\mathcal{T}^{d}_{t})$ to $(X,\mathcal{T}^{d^{\prime}}_{t+\varepsilon})$. The second map is obtained symmetrically. ∎
4.2 Interleaving of Persistence Modules
Applying the functorial constructions of Section 3 yields filtrations of simplicial complexes
$\mathbb{K}^{d},\mathbb{K}^{d^{\prime}}\colon(\mathbb{R},\leq)\longrightarrow\mathbf{Simp}.$
Proposition 9.
For each $n\geq 0$, the persistence modules $H_{n}(\mathbb{K}^{d})$ and $H_{n}(\mathbb{K}^{d^{\prime}})$ are $\varepsilon$-interleaved. No inclusion relations between the complexes are required.
Proof.
The identity maps of the previous lemma induce, functorially, simplicial maps
$K^{d}_{t}\longrightarrow K^{d^{\prime}}_{t+\varepsilon},\qquad K^{d^{\prime}}_{t}\longrightarrow K^{d}_{t+\varepsilon}.$
Functoriality ensures that these maps are compatible with the filtrations and satisfy the coherence conditions required for an $\varepsilon$-interleaving. Applying homology completes the proof. ∎
4.3 Stability of Persistence Diagrams
Combining the above with the Isometry Theorem yields the desired stability result.
Theorem 10 (Stability).
For each $n\geq 0$,
$d_{B}\bigl(\mathsf{Dgm}(H_{n}(\mathbb{K}^{d})),\mathsf{Dgm}(H_{n}(\mathbb{K}^{d^{\prime}}))\bigr)\leq\varepsilon.$
Proof.
By the previous proposition, the persistence modules are $\varepsilon$-interleaved. The result follows from the Isometry Theorem (Theorem 1). ∎